# WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs

### v1

@@ 3. Overview and an Example
If there is a satisfying assignment of the permission variables with permission tokens, then intuitively, the operators in the dataflow program are consistent in memory accesses, and there will not be any data races. In Section 5, we show that this implies that any possible schedule will converge to the canonical schedule we use. For the example in Figure 3, a satisfying assignment is $p_{2}=p_{7}=p_{15}=p_{16}=p_{17}=p_{18}=\texttt{write}$ with all other variables set to 00. This assignment essentially passes the write permission across the loop of operators 2, 5, 7, 8, 9.
4. Bisimulation on the Canonical Schedule
@@
Now, we detail the first half of our translation validation technique, which proves via cut-bisimulation that the LLVM program is equivalent to the dataflow program when restricted to a canonical schedule.
@@
Given a dataflow program and LLVM program, their semantics are defined by transition systems $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$, respectively, where $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{llvm}}$ are the sets of initial configurations and $\mathcal{C}_{\mathrm{df}}$ and $\mathcal{C}_{\mathrm{llvm}}$ are the sets of all configurations.
@@
We use $\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})$ to denote the state of the memory in a dataflow configuration $c_{\mathrm{df}}$, and similarly $\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})$ for the memory in an LLVM configuration $c_{\mathrm{llvm}}$. We will omit the subscripts when it is clear from the context. For the dataflow semantics, we use $\to^{o}_{\mathrm{df}}$ to denote the (deterministic) firing of a particular operator $o$. For a transition relation $\to$, we use $\to^{*}$ for the reflexive-transitive closure of $\to$, and $\to^{+}$ for the transitive closure of $\to$. A configuration $c$ is terminating/final if there exists no configuration $c^{\prime}$ such that $c\to c^{\prime}$.
@@
One important tool that we use is symbolic execution, which is similar to concrete execution but performed with symbolic expressions in the configuration. In the following, we establish some notation for symbolic execution. Let $\mathbf{x}$ be a list of variables $\mathbf{x}=(x_{1},\dots,x_{n})$. Let $V$ be the set of all bit-vectors of any lengths. We use $V(\mathbf{x})$ to denote the set of symbolic expressions constructed from constants in $V$, variables in $\mathbf{x}$, and any bit-vector operation we could perform in our semantics (e.g. 32-bit zero bit-vector $\texttt{0}_{32}$, or terms like $x_{1}+\texttt{1}_{32}$ are in $V(\mathbf{x})$). For $s\in\{\mathrm{df},\mathrm{llvm}\}$, we use $\mathcal{C}_{s}(\mathbf{x})$ to represent the set of symbolic configurations with free variables $\mathbf{x}$ in the semantics $s$. A symbolic configuration $c(\mathbf{x})\in\mathcal{C}_{s}(\mathbf{x})$ is similar to a configuration but over symbolic expressions $V(\mathbf{x})$ instead of only $V$ (for example, a symbolic dataflow configuration can have symbolic expressions such as $x_{1}+\texttt{1}_{32}$ in the channels; a symbolic LLVM configuration can have symbolic expressions stored in variables as opposed to only concrete values in $V$). If we have a substitution $\sigma:\{x_{1},\dots,x_{n}\}\to V(\mathbf{y})$ from $\mathbf{x}$ to symbolic expressions over some other list of variables $\mathbf{y}$, we denote $c(\sigma)\in\mathcal{C}_{s}(\mathbf{y})$ to be the result of the substitution. We also lift the transition relation $\to_{s}$ (for $s\in\{\mathrm{df},\mathrm{llvm}\}$) to symbolic configurations: we use $c(\mathbf{x})\to_{s}c^{\prime}(\mathbf{x})\mid\varphi(\mathbf{x})$ to denote the symbolic execution from $c(\mathbf{x})$ to $c^{\prime}(\mathbf{x})$ with path condition $\varphi(\mathbf{x})$.
The dataflow transition relation $\to_{\mathrm{df}}$ could be nondeterministic (a configuration can have multiple outgoing transitions), while $\to_{\mathrm{llvm}}$ is deterministic. However, for this section, we only care about a canonical schedule for the execution of operators. We denote the transition relation restricted to this canonical schedule by $\to_{\mathrm{df\mid canon}}\subseteq\to_{\mathrm{df}}$, and we focus on this deterministic transition for this section. The exact meaning of this canonical schedule will be explained in Section 4.2, but essentially, it executes dataflow operators in the same order their corresponding LLVM instructions would execute.
The equivalence property that we want to check is that for any pair of concrete initial states $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$,
  • $c_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for a final state $c^{\prime}_{\mathrm{df}}$ iff $c_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for a final state $c^{\prime}_{\mathrm{llvm}}$, in which case $\mathrm{mem}(c^{\prime}_{\mathrm{df}})=\mathrm{mem}(c^{\prime}_{\mathrm{llvm}})$.
  • $c_{\mathrm{df}}$ is non-terminating iff $c_{\mathrm{llvm}}$ is non-terminating, and there are infinitely many pairs of states in their traces with equal memory.
@@
Since the dataflow and LLVM traces could both be unbounded, naive symbolic execution from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$ may not terminate. So we need to prove a coinductive invariant about the pairs of traces starting from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$, and such an invariant is a cut-bisimulation relation.
Definition 0 (Cut-bisimulation).
@@
Let $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ be a binary relation. We say that $\mathcal{R}$ is a cut-bisimulation iff for any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$, if $c_{\mathrm{df}}\to_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for some $c^{\prime}_{\mathrm{df}}$, then there are $(c^{\prime\prime}_{\mathrm{df}},c^{\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime\prime}_{\mathrm{df}}$ and $c_{\mathrm{llvm}}\to^{+}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$. Conversely, if $c_{\mathrm{llvm}}\to_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for some $c^{\prime}_{\mathrm{llvm}}$, then there are $(c^{\prime}_{\mathrm{df}},c^{\prime\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime\prime}_{\mathrm{llvm}}$ and $c_{\mathrm{df}}\to^{+}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$.
This is a weaker variant of the usual notion of bisimulation where we only allow one step before synchronizing instead of allowing one or more steps. Work by Kasampalis et al. (keq, ) provides more detail about this variant of bisimulation.
@@
To check equivalence, we look for cut-bisimulations that are memory-synchronizing:
Definition 0 ().
@@
A binary relation $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ is memory-synchronizing iff
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\subset\mathcal{R}$, $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$.
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})$ with $c_{\mathrm{df}},c_{\mathrm{llvm}}$ final and $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
@@
If there exists a cut-bisimulation $\mathcal{R}$ that is memory synchronizing, then we can conclude via a coinductive argument that the two programs are equivalent.
In the following subsections, we first describe how we represent a cut-bisimulation relation via cut points (Section 4.1), and how to automatically check that a proposed relation is actually a cut-bisimulation via symbolic execution (Section 4.2). Then in Section 4.3, we describe how the cut points are inferred using hints from the compiler.
4.1. Describing a Cut-Bisimulation via Cut Points

### v2

@@ 3. Overview and an Example
If there is a satisfying assignment of the permission variables with permission tokens, then intuitively, the operators in the dataflow program are consistent in memory accesses, and there will not be any data races. In Section 5, we show that this implies that any possible schedule will converge to the canonical schedule we use. For the example in Figure 3, a satisfying assignment is $p_{2}=p_{7}=p_{15}=p_{16}=p_{17}=p_{18}=\texttt{write}$ with all other variables set to 00. This assignment essentially passes the write permission across the loop of operators $2\to 5\to 7\to 8\to 9\to 2$.
4. Equivalence on a Canonical Schedule via Cut-Simulation
@@
We detail the first half of our translation validation technique, which proves via cut-simulation that the LLVM program is equivalent to the dataflow program when restricted to a canonical schedule of dataflow operators.
Most of our formulation is independent of the source language. Hence, instead of using LLVM directly, we assume some sequential and imperative input language called Seq such that 1) Seq is deterministic; 2) Seq has a sequentially consistent memory model (as do dataflow programs); 3) Seq programs operate on fixed-width integer values (as do dataflow programs).
@@
Throughout this section, we fix a dataflow program and a Seq program, and we denote their operational semantics as transition systems $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$, respectively, where $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{seq}}$ are the sets of initial configurations, and $\mathcal{C}_{\mathrm{df}}$ and $\mathcal{C}_{\mathrm{seq}}$ are the sets of all configurations.
We use $\mathsf{Mem}_{\mathrm{df}}(c_{\mathrm{df}})$ to denote the state of the memory in a dataflow configuration $c_{\mathrm{df}}$, and similarly $\mathsf{Mem}_{\mathrm{seq}}(c_{\mathrm{seq}})$ for the memory in a Seq configuration $c_{\mathrm{seq}}$. We will omit the subscripts when it is clear from the context. We assume that the memory map is modeled in the same way in both semantics, so that we can directly compare them by $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$. Additionally, both the dataflow and Seq programs have input parameters (e.g., the pointer %A in Figure 2(b)); thus, we assume a relation $\sim_{\mathcal{I}}\ \subseteq\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{seq}}$ on initial configurations that guarantees equal values on all input parameters.
@@
For a transition relation $\to\ \in\{\to_{\mathrm{df}},\to_{\mathrm{seq}}\}$, we use $\to^{*}$ for the reflexive-transitive closure of $\to$, and $\to^{+}$ for the transitive closure of $\to$. A configuration $c$ is terminating/final if there exists no configuration $c^{\prime}$ such that $c\to c^{\prime}$. A trace is a finite or infinite sequence of configurations $\tau=c_{1}\dots c_{i}\dots$ such that $c_{i}\to c_{i+1}$ for any $i$. A trace $\tau$ is complete if $\tau$ is infinite or the last configuration of $\tau$ is final. We define $\mathsf{Traces}_{\mathrm{df}}$ and $\mathsf{Traces}_{\mathrm{seq}}$ to be the set of all traces in the dataflow and Seq semantics, respectively. For the dataflow semantics, we use $\to^{o}_{\mathrm{df}}$ to denote the deterministic firing of a particular operator $o$.
@@
One important tool that we use is symbolic execution, and in the following, we establish some notation for symbolic execution. Let $\mathbf{x}$ be a list of variables $\mathbf{x}=(x_{1},\dots,x_{n})$. Let $V$ be the set of all bit-vectors of a predetermined length (for modeling fixed-width integers). We use $\mathsf{Terms}(\mathbf{x})$ to denote the set of symbolic expressions constructed from constants in $V$, variables in $\mathbf{x}$, and any bit-vector operation we could perform in our semantics (e.g., a 32-bit zero bit-vector $\texttt{0}_{32}$, or a term like $x_{1}+\texttt{1}_{32}$ are in $\mathsf{Terms}(\mathbf{x})$). For $s\in\{\mathrm{df},\mathrm{seq}\}$, we use $\mathcal{C}_{s}(\mathbf{x})$ to represent the set of symbolic configurations with free variables $\mathbf{x}$ in the semantics $s$. A symbolic configuration $c(\mathbf{x})\in\mathcal{C}_{s}(\mathbf{x})$ is similar to a configuration but over symbolic expressions $\mathsf{Terms}(\mathbf{x})$ instead of only $V$ (for example, a symbolic dataflow configuration can have symbolic expressions such as $x_{1}+\texttt{1}_{32}$ in the channels). If we have a substitution $\sigma:\{x_{1},\dots,x_{n}\}\to\mathsf{Terms}(\mathbf{y})$ from $\mathbf{x}$ to symbolic expressions over some other list of variables $\mathbf{y}$, we denote $c(\sigma)\in\mathcal{C}_{s}(\mathbf{y})$ to be the result of the substitution. We use $c(\mathbf{x})\to_{s}c^{\prime}(\mathbf{x})\mid\varphi(\mathbf{x})$ to denote symbolic execution from $c(\mathbf{x})$ to $c^{\prime}(\mathbf{x})$ with path condition $\varphi(\mathbf{x})$.
A dataflow program does not have a notion of a return value, and the only observable state after finishing execution is the modified global memory. Thus we first define when two traces are equivalent based on their memory state.
Definition 0 (Memory-synchronizing traces).
Let $\tau_{\mathrm{df}}=c^{1}_{\mathrm{df}}\dots c^{i}_{\mathrm{df}}\dots$ and $\tau_{\mathrm{seq}}=c^{1}_{\mathrm{seq}}\dots c^{j}_{\mathrm{seq}}\dots$ be two complete traces starting from $c^{1}_{\mathrm{df}}$ and $c^{1}_{\mathrm{seq}}$, respectively. The traces $\tau_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing iff $\mathsf{Mem}(c^{1}_{\mathrm{df}})=\mathsf{Mem}(c^{1}_{\mathrm{seq}})$, and
  • If $\tau_{\mathrm{df}}$ is finite, then $\tau_{\mathrm{seq}}$ is finite and $\mathsf{Mem}(c^{\prime}_{\mathrm{df}})=\mathsf{Mem}(c^{\prime}_{\mathrm{seq}})$, where $c^{\prime}_{\mathrm{df}}$ and $c^{\prime}_{\mathrm{seq}}$ are the final states of $\tau_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$, respectively.
  • If $\tau_{\mathrm{df}}$ is infinite, then $\tau_{\mathrm{seq}}$ is infinite, and for any $n\in\mathbb{N}$, there are $i,j\geq n$ such that $\mathsf{Mem}(c^{i}_{\mathrm{df}})=\mathsf{Mem}(c^{j}_{\mathrm{seq}})$.
In other words, two complete dataflow and Seq traces are memory-synchronizing if they are both finite and have the same memory state in the initial and final configurations (i.e., both programs terminate and have the same final memory state); or they are both infinite and they synchronize in the same memory state infinitely often.
Now, we may define program equivalence, which describes when two matching initial configurations for dataflow and Seq exhibit the same observable behaviors:
Definition 0 (Program equivalence).
Two programs $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$ are equivalent if for any pair of initial states $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, if we take any complete dataflow trace $\tau_{\mathrm{df}}=c_{\mathrm{df}}\dots$ and the unique complete Seq trace $\tau_{\mathrm{seq}}=c_{\mathrm{seq}}\dots$ (unique since Seq is deterministic), then $\tau_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing.
Intuitively, the dataflow program and Seq program are equivalent if any pair of complete traces from corresponding initial states are memory-synchronizing; i.e., they either both terminate with the same memory state, or both run forever and synchronize infinitely often.
Our final goal is to achieve Definition 2. We will do so by first verifying a weaker property of equivalence along a particular, deterministic canonical schedule of the dataflow program. Then, in Section 5, we show that this canonical schedule suffices to prove equivalence for all possible dataflow schedules.
First, we introduce canonical schedules, which determinize dataflow programs using the information contained in Seq traces:
Definition 0 (Canonical schedule).
A canonical schedule is a function $\mathsf{Canon}:\mathcal{C}_{\mathrm{df}}\times\mathsf{Traces}_{\mathrm{seq}}\to\mathsf{Traces}_{\mathrm{df}}$ that selects a dataflow trace using a Seq trace, such that
  • For any $c_{\mathrm{df}}\in\mathcal{C}_{\mathrm{df}}$ and any non-empty Seq trace $\tau_{\mathrm{seq}}$, $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is a dataflow trace starting in $c_{\mathrm{df}}$.
  • If $\tau_{\mathrm{seq}}$ is finite, then $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is finite.
  • For any $c_{\mathrm{df}}\in\mathcal{C}_{\mathrm{df}}$, if $\tau^{\prime}_{\mathrm{seq}}$ is a prefix of $\tau_{\mathrm{seq}}$, then $\mathsf{Canon}(c_{\mathrm{df}},\tau^{\prime}_{\mathrm{seq}})$ is a prefix of $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$.
We then specialize program equivalence (Definition 2) by restricting our attention to those canonical dataflow schedules:
Definition 0 (Program equivalence on a canonical schedule).
Let $\mathsf{Canon}$ be a canonical schedule. Two programs $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$ are equivalent on the canonical schedule $\mathsf{Canon}$ if for any pair of initial states $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, if we take the unique complete Seq trace $\tau_{\mathrm{seq}}=c_{\mathrm{seq}}\dots$, then $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is a complete dataflow trace, and $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing.
That is, instead of enforcing that all complete dataflow traces are memory-synchronizing to the complete Seq trace, we only require that there exists one such trace selected by some canonical schedule $\mathsf{Canon}$. Later in Section 5, we separately verify the confluence property (Definition 1), so that Definition 1 and Definition 4 together imply Definition 2 (Proposition 2).
Canonical Schedules for RipTide
[…]
$\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{llvm}}):=\begin{cases}\epsilon&\text{If }\tau_{\mathrm{llvm}}=\epsilon\\ \tau_{\mathrm{df}}\cdot\mathsf{Canon}(c^{\prime}_{\mathrm{df}},\tau^{\prime}_{\mathrm{llvm}})&\text{If }\tau_{\mathrm{llvm}}=c_{\mathrm{llvm}}\cdot\tau^{\prime}_{\mathrm{llvm}}\end{cases}$
[…]
In the following sections, we outline the steps to verify that the dataflow and Seq programs satisfy Definition 4. In Section 4.1, we first describe our general method of using a coinductive invariant called cut-simulation that implies the property in Definition 4. In Section 4.2, we show how cut-simulation can be symbolically represented by cut points. In Section 4.3, we present the algorithm to check that a set of cut points does form a cut-simulation. Finally in Section 4.4, we describe our heuristics to infer cut points for the compilation from LLVM to dataflow programs.
4.1. Proving Equivalence on the Canonical Schedule via Cut-Simulation
@@
Traces from $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{seq}}$ are unbounded in general, so to prove the equivalence on a canonical schedule (Definition 4), we need to verify a coinductive invariant about the pairs of traces starting from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{seq}})$, and such an invariant is a cut-simulation relation (Kasampalis et al., 2021).
Definition 0 (Cut-simulation).
@@
Let $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}$ be a binary relation. We say that $\mathcal{R}$ is a cut-simulation between $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$ and $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ iff for any $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$, if $c_{\mathrm{seq}}\to_{\mathrm{seq}}c^{\prime}_{\mathrm{seq}}$ for some $c^{\prime}_{\mathrm{seq}}$, then there are $(c^{\prime}_{\mathrm{df}},c^{\prime\prime}_{\mathrm{seq}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{seq}}\to^{*}_{\mathrm{seq}}c^{\prime\prime}_{\mathrm{seq}}$ and $c_{\mathrm{df}}\to^{+}_{\mathrm{df}}c^{\prime}_{\mathrm{df}}$.
Intuitively, a cut-simulation states that if the two transition systems are synchronized in $\mathcal{R}$, then if we can take a step on the Seq side, we can step the two transition systems further to re-synchronize them in $\mathcal{R}$.
Note that in our setting, we are verifying that the dataflow program simulates the Seq program (i.e., the target program simulates the source program), instead of the other way around which is common in compiler verification work. This is because our target dataflow program has more nondeterministic behavior than the source Seq program, whereas in many other cases, the source program has more behavior (such as undefined behavior in C) than the concrete target program.
@@
To satisfy our desired equivalence properties, we look for cut-simulations that are memory-synchronizing:
Definition 0.
@@
A binary relation $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}$ is memory-synchronizing iff
  • For any $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$, $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$.
  • For any initial configurations $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$.
  • For any $(c_{\mathrm{df}},c_{\mathrm{seq}})$ with $c_{\mathrm{df}},c_{\mathrm{seq}}$ final and $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$.
The second and third conditions enforce that $\mathcal{R}$ should include all pairs of corresponding initial states, as well as final states; thus $\mathcal{R}$ covers all complete traces.
@@
If there exists a memory-synchronizing cut-simulation $\mathcal{R}$, then we can conclude via a coinductive argument that the two programs are equivalent on a canonical schedule (Definition 4).
4.2. Describing a Cut-Simulation via Cut Points